OpenStudy (anonymous):

f(x)=2X+1 and g(x)=3x-2 find (f-g)(4)

jimthompson5910 (jim_thompson5910):

(f-g)(4) is the same as f(4) - g(4)

jimthompson5910 (jim_thompson5910):

so you need to evaluate f(4) and g(4), then subtract

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so I plug in the 4 where there are x and then solve?

jimthompson5910 (jim_thompson5910):

f(x)=2x+1 f(4)=2(4)+1 f(4) = ???

OpenStudy (anonymous):

thank you . I have it now

jimthompson5910 (jim_thompson5910):

ok great